OpenStudy (anonymous):

A piece of green jade has amass of 26.123g. if the sample of jade displaces 50.0mL of water to 57.5mL in a graduated cylinder, what is the density of the jade?

OpenStudy (nottim):

density formula= p=m/v

OpenStudy (nottim):

So what next?

OpenStudy (anonymous):

is it 26.123/50.0

OpenStudy (nottim):

Almost. You got the volume off

OpenStudy (nottim):

You have to do something extra with the water

OpenStudy (anonymous):

not sure if this is correct 57.5-50.0 = the volume

OpenStudy (nottim):

That's what I thought

OpenStudy (anonymous):

ok so 57.5-50.0= 7.5 26.123/7.5=3.48g/mL

OpenStudy (nottim):

I can't tell you the answer, dearest. But your steps look good.

OpenStudy (anonymous):

would you mind helping me with another one

OpenStudy (nottim):

sorry, im a bit busy!

OpenStudy (anonymous):

its ok thank you